OpenStudy (anonymous):

What is the value of s in the equation 3r = 10 + 5s, when r = 10? 4 8 100 200

OpenStudy (anonymous):

3r = 10 + 5s Plug in 10 for 'r': 3(10) = 10 + 5s What's 3 * 10?

OpenStudy (anonymous):

substitute 10 for r.

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

@Rafyy Can you multiply 3 x 10?

OpenStudy (anonymous):

this gets you to 3*10=10+5s .

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Yes, then can you subtract 30 - 10?

OpenStudy (anonymous):

Yes, so we have: 30 = 10 + 5s Now subtract 10 to both sides, what's 30 - 10/

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Yes, so we have 20 = 5s What's 20 / 5?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Divide 20 \(\div\) 5 now. :)

OpenStudy (anonymous):

4!

OpenStudy (anonymous):

Yes, that's our answer.

OpenStudy (anonymous):

now you subtract 10 from both sides to cancel things out and to help isolate the variable.

OpenStudy (anonymous):

lol i type so slow

OpenStudy (anonymous):

THANKS GUYS!! =D